# Benchmarking Pedestrian Dynamics Models for Common Scenarios: An Evaluation of Force-Based Models

### v1

@@ Abstract
Abstract: Extensive research in pedestrian dynamics has primarily focused on crowded conditions and associated phenomena, such as lane formation, evacuation, etc. Several force-based models have been developed to predict the behavior in these situations. In contrast, there is a notable gap in terms of investigations of the moderate-to-low density situations. These scenarios are extremely commonplace across the world, including the highly populated nations like India. Additionally, the details of force-based models are expected to show significant effects at these densities, whereas the crowded, nearly packed, conditions may be expected to be governed largely by contact forces. In this study, we address this gap and comprehensively evaluate the performance of different force-based models in some common scenarios. Towards this, we perform controlled experiments in four situations: avoiding a stationary obstacle, position-swapping by walking toward each other, overtaking to reach a common goal, and navigating through a maze of obstacles. The performance evaluation consists of two stages and six evaluating parameters - successful trajectories, overlapping proportion, oscillation strength, path smoothness, speed deviation, and travel time. Firstly, models must meet an eligibility criterion of at least 80% successful trajectories and secondly, the models are scored based on the cutoff values established from the experimental data. We evaluated five force-based models where the best one scored 57.14%. Thus, our findings reveal significant shortcomings in the ability of these models to yield accurate predictions of pedestrian dynamics in these common situations.
@@ I Introduction
The dynamics of pedestrians has been studied by various models proposed over past few decades by different researchers. These include the perspectives of the motion of individual pedestrians in any given scenario as well as the evolution of the state of groups of pedestrians. The types of models also vary widely and range from force-based, using classical laws of motion karamouzas2014universal ; helbing1995social ; chraibi2010generalized , to heuristic, rule-based approaches moussaid2011simple . These studies have been accompanied by appropriate experimental ones as well, across various nations and situations helbing2007dynamics ; haghani2017following ; appert2018experimental ; moussaid2009experimental . The central problem, one of being able to accurately model pedestrian motion or the state of a group of pedestrians (or a crowd, which usually refers to a large number of pedestrians at high density), is of enormous significance. This is expected to enable the fundamental physics-based design of public infrastructure, where the aspects of safety and mitigation of stampede-like disasters. Such successful models will have a positive impact on public space designs across the world, more significantly in countries like India, with high population densities.
Table 1: Details of experiments (mimicking real-life scenarios) performed with the number of data collected for each case.
Experiment Data Collected
@@
As discussed, various models have been proposed by earlier researchers, showing great promise to mimic different pedestrian scenarios. These need to be thoroughly evaluated so that the correct model can be selected for designing a given scenario. We anticipate that the evaluation of the effectiveness of these models will involve various situations - ranging from those involving the motion of many pedestrians to a few, with or without other obstacles (stationary, may or may not be involving other human beings) around them. The former, or crowded scenarios, include lane formation feliciani2016empirical ; schadschneider2011empirical ; kramer2021social , bottleneck schadschneider2011empirical ; helbing2006analytical , evacuation von2017empirical ; shiwakoti2019review ; heliovaara2012pedestrian ; lui2012constructal ; lui2013constructal , and crowd intersection gorrini2013empirical ; shiwakoti2015empirical ; lian2015experimental , extensively documented in the literature. Evaluation parameters for these scenarios include fundamental diagrams seyfried2005fundamental ; seyfried2007fundamental ; seyfried2010enhanced , density measure lerner2009data , velocity fields liddle2022microscopic , etc. There are empirical studies on different crowded scenarios in the literature haghani2018crowd ; haghani2020empirical . These studies provide benchmarks for assessing model performance in crowded conditions. However, note that all these situations involve a high density crowd, at closely packed conditions. Under such conditions, the close contact interactions are expected to govern pedestrian motion rather than the fundamental interactions between them, which might be active when they have some gaps around each other to move and manoeuvre. Thus, almost all reasonable force-based models (and maybe even rule based ones) may be expected to converge to similar answers for such highly packed scenarios. Interesting differences may be anticipated in moderate-to-low density situations, where the details of the interaction between pedestrians at some distance (i.e. not in close contact) are likely to influence the motion of individual pedestrians. Interestingly, in sharp contrast to crowded scenarios, much fewer studies focus on moderate-to-low density scenarios. Typically, such studies focus on the measurement of specific features like personal space shan2014critical ; parisi2016experimental , gait features cao2018stepping ; renaudin2012step ; racic2009experimental , etc., rather than the comparison of motion of pedestrians in situations from experiments and models. Thus, since such situations are not clearly established, the aspect of model evaluation with varying details of forces (and rules) remains incomplete. Hence, there exists a critical gap in terms of establishing benchmarks for evaluating pedestrian dynamics models, which we would address in this article. Note, a gap also exists for such comparisons involving Indian pedestrians, as most of the earlier investigations were performed in western or eastern nations. In this context, we note that such moderate-to-low density conditions are extremely common in various public spaces at all times, across India and other nations with large population densities. Our investigations here also extends to high density situations involving one moving pedestrian among many stationary obstacles (may or may not be humans). In this context, in a previous study, we have shown that the behavior of a single pedestrian around a non-living obstacle is similar to that around a similar-sized human being jainAutomatic . Ideally, we expect all the models to perform reasonably well at both high, moderate and low densities, but our study here proves the contrary (as shown later in the article).
@@
From the perspective discussed above, this study addresses this gap for various force-based models in the literature. We focus on common situations at the moderate-to-low density regimes, where the details of interactions will govern the pedestrian motion. In the process, we have two broad objectives. Firstly, the benchmarking criteria for such situations are established, such that any future models can be evaluated similarly. Secondly, various force-based models present in the literature are evaluated within this criteria. Towards this, our study performs controlled experiments across four commonplace real-life scenarios (Figure 1), and respective data collection is given in Table 1. As explained later, these situations include one or two pedestrians, in the absence or presence of other stationary obstacles. These obstacles may represent other human beings (stationary) or inanimate objects (like pillars or similar objects). The number of obstacles is varied to traverse various regimes of density. Bottinelli and Silverburg observe instances where high-density crowds behave similarly to soft solids bottinelli2019dense . Here, the term “density” is used in the broader sense of the total number of objects within the experimental area, including the pedestrians. For evaluation across models, we use six evaluating parameters - successful trajectories, overlapping proportion, oscillation strength, path smoothness, speed deviation, and travel time, as explained in the article. There are many force-based models present in the literature karamouzas2014universal ; helbing2000simulating ; helbing1995social ; chraibi2009quantitative ; chraibi2010generalized ; hu2023anticipation . We evaluate five that are frequently used for applications - Universal Power Law (UPL) karamouzas2014universal , Social Force Model - circular (SFMc) helbing2000simulating , Social Force Model - elliptical (SFMe) helbing1995social , Centrifugal Force Model - circular (CFMc) chraibi2009quantitative , and Centrifugal Force Model - elliptical (CFMe) chraibi2010generalized . These models were tested against the selected situations and evaluated using a scoring system based on a two-stage evaluation process, as explained later. In the first stage, models are assessed against specific eligibility criteria to determine their suitability. In the second stage, a finer comparison of the models is performed based on cutoff values established from experimental data. As highlighted later, our findings reveal significant deficiencies in the ability of force-based models to predict experimental pedestrian behavior. To regenerate the experimental scenarios and calculate evaluating parameters for scoring, initial position, final position, and desired speed for each volunteer and maximum travel time in each scenario are provided https://github.com/kanika201293/Pedestrian-Experimental-Data.
@@ II Methodology
Figure 2: Schematic of all the experimental setups within a boundary defined as $x=0\;m$ to $x=10\;m$ and $y=-1.75\;m$ to $y=1.75\;m$. The measurement region spans from $x=2\;m$ to $x=8\;m$. For future reference, volunteer/simulated pedestrian A is represented by a filled circle, and an empty circle represents volunteer/simulated pedestrian B.
@@ II.1 Experimental Setup
In this study, we conducted controlled experiments for four different low to moderate density real-life-based scenarios. Note, all these scenarios involve 1-2 pedestrians, but some also have a high density of stationary obstacles. However, all of these represent commonly occurring everyday situations. Note, from the point of view of the moving pedestrian, an obstacle is equivalent to a stationary similar-sized human, as shown in our earlier study jainAutomatic . We selected obstacles that have similar diameter as a human being. Table 1 provides details of the experiments and data collection procedure. The experimental boundary was defined as a rectangular area with $x=0\;m$ to $x=10\;m$ and $y=-1.75\;m$ to $y=1.75\;m$, with the measurement region spanning from $x=2\;m$ to $x=8\;m$. The experiments conducted are as follows:
@@
II.2 Evaluating Parameters
The following are the parameters used in this study to evaluate the performance of any model, for any given scenario.
@@ II.2 Evaluating Parameters
Overlapping Proportion: “Overlapping” between two pedestrians or between a pedestrian and an obstacle occurs when their geometrical form overlaps. In this study, all the pedestrians and obstacles are modeled as circles. A measure of the extent of overlap is defined earlier by Chraibi et. al. chraibi2010generalized , as follows:
@@
Here, $N$ represents the total count of simulated pedestrians, $A_{i}$ and $A_{j}$ denote the areas of the circular pedestrians $i$ and $j$, respectively, and the variable $A_{ij}$ represents the overlapping area between the two. In (1), the expression is normalized by the variable $nov$, representing the number of simulation steps where overlapping appears. $nov$ is 00 if there is no overlapping during the whole simulation. In order to evaluate the models thoroughly, the maximum overlap is also calculated in some cases.
Oscillation Strength: Backward movements of a pedestrian (opposite to the direction of the goal) during a simulation are called oscillations. Such movements are unrealistic in nature and need to be addressed in a model. To quantify the oscillations, we used the following measure introduced by Chraibi et al. chraibi2010generalized :
@@
Here, $N$ represents the total count of simulated pedestrians, $\overrightarrow{v_{i}}$ is the current velocity, and $\overrightarrow{v_{i}^{0}}$ is the desired velocity of pedestrian $i$. In (2), the expression is normalized by the variable $nos$, representing the number of simulation steps where oscillation appears; $nos$ is set to 00 if there is no oscillation during the whole simulation. Note that the desired velocity is the desired speed of the pedestrian towards the goal.
Path Smoothness: This parameter measures the smoothness of a path by identifying any sudden changes in path direction. It is calculated by measuring the maximum absolute change in the path angle which is calculated at each time instant. Let $\overrightarrow{P_{t}}$ be the position vector of the pedestrian at time instant $t$. The displacement vector $\overrightarrow{V_{t}}$ is defined as the change between position vectors $\overrightarrow{P_{t}}$ and $\overrightarrow{P_{t+1}}$. The angular deviation of path, $\theta_{t}$ is the angle between the displacement vectors $\overrightarrow{V_{t}}$ and $\overrightarrow{V_{t+1}}$. Path smoothness, $\theta_{i}$, for pedestrian $i$, is then the maximum of the absolute value of angles $\theta_{t}$. The lower value of $\theta_{i}$ denotes higher path smoothness for the pedestrian. The quantitative calculation of path smoothness is as follows:
@@
Here, $v_{i}^{0}$ represents the desired speed, and $v_{i}(t)$ represents the speed at time instant $t$, respectively for pedestrian $i$. $nsd$ refers to the number of simulation steps and $N$ refers to the total count of simulated pedestrians.
Travel Time: This parameter represents the total time taken by a pedestrian to traverse the measurement area. During simulation, a trajectory is deemed unsuccessful if the time taken by the simulator exceeds twice the maximum travel time recorded during the experiment.
@@
1) Single values of evaluating parameters for all the scenarios are calculated (shown in Table 2), by averaging over all the trajectories.
2) All the evaluating parameters are calculated in the measurement region to avoid entry/exit effects.
@@ II.3 Experimental Observations
Table 2: Evaluating parameters are calculated from the trajectories for all the experimental setups to define a benchmark to evaluate the functioning of pedestrian dynamics models for low pedestrian density scenarios. The experiments show no unsuccessful trajectories, oscillations, or overlaps. Path smoothness, measured by the change in angle between two consecutive motion vectors, ranges from 0 to 2 degrees. The speed deviation from the desired speed at any instant is not more than 23%. Additionally, the average travel time for volunteers is provided for each case.
@@
SOSP 100% 0.08% 0 $0.74\pm 0.33^{\circ}$ $0.05\pm 0.03$ $4.52\pm 0.89s$
Head-On 100% 0% 0 $0.55\pm 0.12^{\circ}$ $0.02\pm 0.01$ $3.97\pm 0.29s$
Parallel Ped 100% 0% 0 $0.77\pm 0.18^{\circ}$ $0.05\pm 0.04$ $5.21\pm 1.4s$
@@
Case D: 14.96% 100% 0% 0 $0.93\pm 0.36^{\circ}$ $0.09\pm 0.05$ $4.7\pm 0.61s$
In this section, we calculate the evaluating parameters defined in Section II.2 for all experimental setups to establish a benchmark for assessing pedestrian dynamics models in these common scenarios. The values of these parameters are summarized in Table 2. Results show no unsuccessful trajectory in any scenario, which shows that the experimental setups, including all cases of MOSP, are such that: a) each pedestrian is able to find a path to their goal without fail, b) pedestrians are not squeezed or overlapped by more than 50% with other pedestrian or obstacles at any time, c) pedestrians remain within the defined experimental boundaries. Additionally, there are negligible instances of oscillation or overlap observed in any setup during the experiments. Path smoothness was measured by the change in angle between two consecutive motion vectors, and it remained within a low range of 0 to 2 degrees across all setups. This indicates that there were no sudden changes in the direction of motion, even when multiple obstacles were present. The maximum deviation from the desired speed observed among all scenarios was within the range of 23% of the value, demonstrating that pedestrians tend to maintain a consistent speed regardless of obstructions. Furthermore, the average travel time for volunteers was recorded for each case, providing additional insights into pedestrian dynamics. These evaluation parameters serve as a reliable benchmark for testing the effectiveness of pedestrian dynamics models in these situations.

### v2

@@ Abstract
Abstract: Extensive research in pedestrian dynamics has primarily focused on crowded conditions and associated phenomena, such as lane formation, evacuation, etc. Several force-based models have been developed to predict the behavior in these situations. In contrast, there is a notable gap in terms of investigations of the moderate-to-low density situations. These scenarios are extremely commonplace across the world, including the highly populated nations like India. Additionally, the details of force-based models are expected to show significant effects at these densities, whereas the crowded, nearly packed, conditions may be expected to be governed largely by contact forces. In this study, we address this gap and comprehensively evaluate the performance of different force-based models in some common scenarios. Towards this, we perform controlled experiments in four situations: avoiding a stationary obstacle, position-swapping by walking toward each other, overtaking to reach a common goal, and navigating through a maze of obstacles. The performance evaluation consists of two stages and six evaluating metrics - successful trajectories, overlapping proportion, oscillation strength, path smoothness, speed deviation, and travel time. Firstly, models must meet an eligibility criterion of at least 80% successful trajectories and secondly, the models are scored based on the cutoff values established from the experimental data. We evaluated five force-based models where the best one scored 57.14%. Thus, our findings reveal significant shortcomings in the ability of these models to yield accurate predictions of pedestrian dynamics in these common situations.
@@ I Introduction
The dynamics of pedestrians has been studied by various models proposed over past few decades by different researchers. These include the perspectives of the motion of individual pedestrians in any given scenario as well as the evolution of the state of groups of pedestrians. The types of models also vary widely and range from force-based, using classical laws of motion Karamouzas et al. (2014); Helbing and Molnar (1995); Chraibi et al. (2010), to heuristic, rule-based approaches Moussaïd et al. (2011). These studies have been accompanied by appropriate experimental ones as well, across various nations and situations Helbing et al. (2007); Haghani and Sarvi (2017); Appert-Rolland et al. (2018); Moussaïd et al. (2009). The central problem, one of being able to accurately model pedestrian motion or the state of a group of pedestrians (or a crowd, which usually refers to a large number of pedestrians at high density), is of enormous significance. This is expected to enable the fundamental physics-based design of public infrastructure, where the aspects of safety and mitigation of stampede-like disasters. Such successful models will have a positive impact on public space designs across the world, more significantly in countries like India, with high population densities.
Table 1: Details of experiments (mimicking real-life scenarios) performed with the number of runs for each case.
Experiment #\# runs
@@
As discussed, various models have been proposed by earlier researchers, showing great promise to mimic different pedestrian scenarios. These models must be systematically evaluated to determine which one is most appropriate for a given application, based on how well it captures the relevant properties of the scenario. We anticipate that the evaluation of the effectiveness of these models will involve various situations - ranging from those involving the motion of many pedestrians to a few, with or without other obstacles (stationary, may or may not be involving other human beings) around them. The former, or crowded scenarios, include lane formation Feliciani and Nishinari (2016); Schadschneider and Seyfried (2011); Kramer and Wang (2021), bottleneck Schadschneider and Seyfried (2011); Helbing et al. (2006), evacuation Von Krüchten and Schadschneider (2017); Shiwakoti et al. (2019); Heliövaara et al. (2012); Lui et al. (2012, 2013), and crowd intersection Gorrini et al. (2013); Shiwakoti et al. (2015); Lian et al. (2015), extensively documented in the literature. Evaluation metrics for these scenarios include fundamental diagrams Seyfried et al. (2005, 2007, 2010), density measure Lerner et al. (2009), velocity fields Liddle et al. (2022), etc. There are empirical studies on different crowded scenarios in the literature Haghani and Sarvi (2018); Haghani (2020). These studies provide benchmarks for assessing model performance in crowded conditions. However, note that all these situations involve a high density crowd, at closely packed conditions. Under such conditions, the close contact interactions are expected to govern pedestrian motion rather than the fundamental interactions between them, which might be active when they have some gaps around each other to move and manoeuvre. Thus, almost all reasonable force-based models (and maybe even rule based ones) may be expected to converge to similar answers for such highly packed scenarios. Interesting differences may be anticipated in moderate-to-low density situations, where the details of the interaction between pedestrians at some distance (i.e. not in close contact) are likely to influence the motion of individual pedestrians. Interestingly, in sharp contrast to crowded scenarios, much fewer studies focus on moderate-to-low density scenarios. Typically, such studies focus on the measurement of specific features like personal space Shan et al. (2014); Parisi et al. (2016), gait features Cao et al. (2018); Renaudin et al. (2012); Racic et al. (2009), etc., rather than the comparison of motion of pedestrians in situations from experiments and models. Thus, since such situations are not clearly established, the aspect of model evaluation with varying details of forces (and rules) remains incomplete. Hence, there exists a critical gap in terms of establishing benchmarks for evaluating pedestrian dynamics models, which we would address in this article. Note, a gap also exists for such comparisons involving Indian pedestrians, as most of the earlier investigations were performed in western or eastern nations. In this context, we note that such moderate-to-low density conditions are extremely common in various public spaces at all times, across India and other nations with large population densities. Our investigations here also extends to high density situations involving one moving pedestrian among many stationary obstacles (may or may not be humans). In this context, in a previous study, we have shown that the behavior of a single pedestrian around a non-living obstacle is similar to that around a similar-sized human being Jain et al. . Ideally, we expect all the models to perform reasonably well at both high, moderate and low densities, but our study here proves the contrary (as shown later in the article).
@@
From the perspective discussed above, this study addresses this gap for various force-based models in the literature. We focus on common situations at the moderate-to-low density regimes, where the details of interactions will govern the pedestrian motion. In the process, we have two broad objectives. Firstly, the benchmarking criteria for such situations are established, such that any future models can be evaluated similarly. Secondly, various force-based models present in the literature are evaluated within this criteria. Towards this, our study performs controlled experiments across four commonplace real-life scenarios (Figure 1), and respective data collection is given in Table 1. As explained later, these situations include one or two pedestrians, in the absence or presence of other stationary obstacles. These obstacles may represent other human beings (stationary) or inanimate objects (like pillars or similar objects). The number of obstacles is varied to traverse various regimes of density. Bottinelli and Silverburg observe instances where high-density crowds behave similarly to soft solids Bottinelli and Silverberg (2019). Here, the term “density” is used in the broader sense of the total number of objects within the experimental area, including the pedestrians. For evaluation across models, we use six evaluating metrics - successful trajectories, overlapping proportion, oscillation strength, path smoothness, speed deviation, and travel time, as explained in the article. There are many force-based models present in the literature Karamouzas et al. (2014); Helbing et al. (2000); Helbing and Molnar (1995); Chraibi et al. (2009, 2010); Hu et al. (2023). We evaluate five that are frequently used for applications - Universal Power Law (UPL) Karamouzas et al. (2014), Social Force Model - circular (SFMc) Helbing et al. (2000), Social Force Model - elliptical (SFMe) Helbing and Molnar (1995), Centrifugal Force Model - circular (CFMc) Chraibi et al. (2009), and Centrifugal Force Model - elliptical (CFMe) Chraibi et al. (2010). These models were tested against the selected situations and evaluated using a scoring system based on a two-stage evaluation process, as explained later. In the first stage, models are assessed against specific eligibility criteria to determine their suitability. In the second stage, a finer comparison of the models is performed based on cutoff values established from experimental data. As highlighted later, our findings reveal significant deficiencies in the ability of force-based models to predict experimental pedestrian behavior. To regenerate the experimental scenarios and calculate evaluating metrics for scoring, initial position, final position, and desired speed for each volunteer and maximum travel time in each scenario are provided https://github.com/kanika201293/Pedestrian-Experimental-Data.
@@ II Methodology
Figure 2: Schematic of all the experimental setups within a boundary defined as x=0mx=0\;m to x=10mx=10\;m and y=1.75my=-1.75\;m to y=1.75my=1.75\;m. The measurement region spans from x=2mx=2\;m to x=8mx=8\;m. For future reference, volunteer/simulated pedestrian A is represented by a filled circle, and an empty circle represents volunteer/simulated pedestrian B.
@@ II.1 Experimental Setup
In this study, we conducted controlled experiments for four different low to moderate density real-life-based scenarios. Note, all these scenarios involve 1-2 pedestrians, but some also have a high density of stationary obstacles. However, all of these represent commonly occurring everyday situations. Note, from the point of view of the moving pedestrian, an obstacle is equivalent to a stationary similar-sized human, as shown in our earlier study Jain et al. . We selected obstacles that have similar diameter as a human being. Table 1 provides details of the experiments and data collection procedure. The experimental boundary was defined as a rectangular area with x=0mx=0\;m to x=10mx=10\;m and y=1.75my=-1.75\;m to y=1.75my=1.75\;m, with the measurement region spanning from x=2mx=2\;m to x=8mx=8\;m. The experiments conducted are as follows:
@@
II.2 Evaluating metrics
The following are the metrics used in this study to evaluate the performance of any model, for any given scenario.
@@ II.2 Evaluating metrics
Overlapping Proportion: “Overlapping” between two pedestrians or between a pedestrian and an obstacle occurs when their geometrical form overlaps. In this study, all the pedestrians and obstacles are modeled as circles. A measure of the extent of overlap is defined earlier by Chraibi et. al. Chraibi et al. (2010), as follows:
@@
Here, $N$ represents the total count of simulated pedestrians, $A_{i}$ and $A_{j}$ denote the areas of the circular pedestrians $i$ and $j$, respectively, and the variable $A_{ij}$ represents the overlapping area between the two. In (1), the expression is normalized by the parameter novnov, representing the number of simulation steps where overlapping appears. novnov is 0 if there is no overlapping during the whole simulation. In order to evaluate the models thoroughly, the maximum overlap is also calculated in some cases.
Oscillation Strength: Backward movements of a pedestrian (opposite to the direction of the goal) during a simulation are called oscillations. Such movements are unrealistic in nature and need to be addressed in a model. To quantify the oscillations, we used the following measure introduced by Chraibi et al. Chraibi et al. (2010):
@@
Here, $N$ represents the total count of simulated pedestrians, $\overrightarrow{v_{i}}$ is the current velocity, and $\overrightarrow{v_{i}^{0}}$ is the desired velocity of pedestrian $i$. In (2), the expression is normalized by the parameter nosnos, representing the number of simulation steps where oscillation appears; nosnos is set to 0 if there is no oscillation during the whole simulation. Note that the desired velocity is the desired speed of the pedestrian towards the goal.
Path Smoothness: This metric measures the smoothness of a path by identifying any sudden changes in path direction. It is calculated by measuring the maximum absolute change in the path angle which is calculated at each time instant. Let $\overrightarrow{P_{t}}$ be the position vector of the pedestrian at time instant $t$. The displacement vector $\overrightarrow{V_{t}}$ is defined as the change between position vectors $\overrightarrow{P_{t}}$ and $\overrightarrow{P_{t+1}}$. The angular deviation of path, $\theta_{t}$ is the angle between the displacement vectors $\overrightarrow{V_{t}}$ and $\overrightarrow{V_{t+1}}$. Path smoothness, $\theta_{i}$, for pedestrian $i$, is then the maximum of the absolute value of angles $\theta_{t}$. The lower value of $\theta_{i}$ denotes higher path smoothness for the pedestrian. The quantitative calculation of path smoothness is as follows:
@@
Here, $v_{i}^{0}$ represents the desired speed, and $v_{i}(t)$ represents the speed at time instant $t$, respectively for pedestrian $i$. nsdnsd refers to the number of simulation steps and $N$ refers to the total count of simulated pedestrians.
Travel Time: This metric represents the total time taken by a pedestrian to traverse the measurement area. During simulation, a trajectory is deemed unsuccessful if the time taken by the simulator exceeds twice the maximum travel time recorded during the experiment.
@@
1) Single values of evaluating metrics for all the scenarios are calculated (shown in Table 2), by averaging over all the trajectories.
2) All the evaluating metrics are calculated in the measurement region to avoid entry/exit effects.
@@ II.3 Experimental Observations
Table 2: Evaluating metrics are calculated from the trajectories for all the experimental setups to define a benchmark to evaluate the functioning of pedestrian dynamics models for low pedestrian density scenarios. The experiments show no unsuccessful trajectories, oscillations, or overlaps. Path smoothness, measured by the change in angle between two consecutive motion vectors, ranges from 0 to 2 degrees. The speed deviation from the desired speed at any instant is not more than 23%. Additionally, the average travel time for volunteers is provided for each case.
@@
SOSP 100% 0.08% 0 $74\pm 0.33^{\circ} 0.0$$05\pm 0.03$ $4.52\pm 0.89s$
Head-On 100% 0% 0 $55\pm 0.12^{\circ} 0.0$$02\pm 0.01$ $3.97\pm 0.29s$
Parallel Ped 100% 0% 0 $77\pm 0.18^{\circ} 0.0$$05\pm 0.04$ $5.21\pm 1.4s$
@@
Case D: 14.96% 100% 0% 0 $93\pm 0.36^{\circ} 0.0$$09\pm 0.05$ $4.7\pm 0.61s$
In this section, we calculate the evaluating metrics defined in Section II.2 for all experimental setups to establish a benchmark for assessing pedestrian dynamics models in these common scenarios. The values of these metrics are summarized in Table 2. Results show no unsuccessful trajectory in any scenario, which shows that the experimental setups, including all cases of MOSP, are such that: a) each pedestrian is able to find a path to their goal without fail, b) pedestrians are not squeezed or overlapped by more than 50% with other pedestrian or obstacles at any time, c) pedestrians remain within the defined experimental boundaries. Additionally, there are negligible instances of oscillation or overlap observed in any setup during the experiments. Path smoothness was measured by the change in angle between two consecutive motion vectors, and it remained within a low range of 0 to 2 degrees across all setups. This indicates that there were no sudden changes in the direction of motion, even when multiple obstacles were present. The maximum deviation from the desired speed observed among all scenarios was within the range of 24% of the value, demonstrating that pedestrians tend to maintain a consistent speed regardless of obstructions. Furthermore, the average travel time for volunteers was recorded for each case, providing additional insights into pedestrian dynamics. These evaluation metrics serve as a reliable benchmark for testing the effectiveness of pedestrian dynamics models in these situations.